### v3

@@ Appendix A Institution User Quotations on Evaluation
Appendix B Linear vs Logarithmic Fits
For readers who want the formal evidence behind Fig. 3: we fit both a straight line and a logarithmic curve to the same elicited points and compared their goodness-of-fit using standard metrics. Smaller is better for SSE (sum of squared errors) and RMSE (root-mean-square error); larger is better for $R^{2}$ (coefficient of determination). AIC (Akaike Information Criterion) and BIC (Bayesian Information Criterion) are model-selection scores that penalize extra parameters; lower is better.
@@ Appendix B Linear vs Logarithmic Fits
$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}},\qquad\mathrm{AIC}=n\ln(\mathrm{SSE}/n)+2k,\quad\mathrm{BIC}=n\ln(\mathrm{SSE}/n)+k\ln n,$
where $n$ is the number of points and $k$ the number of fitted parameters. As Table 3 shows, the logarithmic model has dramatically smaller SSE/RMSE, much higher $R^{2}$, and much lower AIC/BIC, so its advantage is not an artifact of the extra parameter but a closer match to the elicited judgments. These all confirm, quantitatively, that the log model explains about $95\%$ of the variance ($R^{2}\!=\!0.94$), while the linear model explains very little ($R^{2}\!=\!0.04$), and that the log model’s advantage is not merely due to having one extra parameter.
In other words, the visual impression is also statistically decisive. Below is the quantitative comparison of linear vs. logarithmic fits on the elicited points. To decide whether a straight line or a logarithmic curve better represents the expert data, we fit both forms to the same points and then measure how close each comes. This is what the values in Table 3 show: SSE and RMSE are the remaining gaps between curve and data (smaller is better), $R^{2}$ is the share of variation explained (closer to 1 is better), and AIC/BIC compare models while penalizing complexity (lower is better). Even after each curve is fitted as well as possible to the data, the log model has dramatically smaller errors, much higher $R^{2}$, and far lower AIC/BIC. This means that the shape of the logarithmic curve matches the human judgments far more closely than the linear line ever can. In other words, it is not just that we chose to use a logarithm; the data themselves show that the logarithmic shape is the right one.
Model $a$ $b$ $c$ SSE RMSE $R^{2}$ AIC BIC
@@
Linear (through origin) cxcx $0.541$ $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
AIC $n\ln(\mathrm{SSE}/n)+2k$, BIC $n\ln(\mathrm{SSE}/n)+k\ln n$; here n=7n{=}7, k=2k{=}2 (log) and k=1k{=}1 (linear). Smaller is better.
Table 3: Goodness-of-fit statistics for the user 1 minor-error dataset ($x$ in pages; $E$ in minor points). Lower SSE/RMSE/AIC/BIC and higher $R^{2}$ indicate better fit. Data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8).
The standard indicators reported in Table 3 show how well a curve follows the observed data points. They illustrate why the logarithmic model offers a markedly superior fit than the linear one. [^12] SSE (sum of squared errors): the total squared distance between the curve and the data (smaller is better). RMSE: the average size of that distance (smaller means the model sits almost on top of the data). $R^{2}$: the fraction of variability explained by the curve ($R^{2}\approx 1$ means excellent, $R^{2}\approx 0$ means poor). AIC and BIC: compact comparison measures that reward better fit but penalize extra parameters (lower is better). Thus, even without going into formulas, the table makes visible that the non-linear model is overwhelmingly better at reflecting real human tolerance. When multiple elicited points are available, we recommend weighted least squares (App. B), down‑weighting sparse or less reliable page sizes. Using the delta method in App. B, we obtain approximate prediction standard errors and plot shaded confidence bands around the fitted curve.

### v4

@@ Appendix A Institution User Quotations on Evaluation
Appendix B Linear vs. Logarithmic Fits
To complement the qualitative trend in Figures 13, we compare linear and logarithmic functional forms by fitting multiple candidate models to the same elicited tolerance points and evaluating goodness-of-fit. We report standard error-based measures (SSE, RMSE), explained variance ($R^{2}$), and information criteria (AIC, BIC), where lower values indicate better fit for SSE/RMSE/AIC/BIC and higher values indicate better fit for $R^{2}$.
Fit metrics.
For observed points $\{(x_{i},y_{i})\}_{i=1}^{n}$ and fitted values $\hat{y}_{i}$, the metrics are defined as:
@@
$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}},\qquad\mathrm{AIC}=n\ln(\mathrm{SSE}/n)+2k,\quad\mathrm{BIC}=n\ln(\mathrm{SSE}/n)+k\ln n,$
where $n$ is the number of points and $k$ is the number of fitted parameters.
Models.
We compare three candidate forms:
(i) Logarithmic: $y=a\ln(1+bx),$
(ii) Linear (with intercept): $y=\alpha+\beta x,$
(iii) Linear (through origin): $y=cx.$
Model (iii) is included because proportional tolerance scaling implicitly assumes a zero intercept (i.e., no tolerance at zero length), whereas Model (ii) provides a standard unconstrained linear baseline.
Table 3 reports fitted parameters and fit statistics for the Institution 1 minor-error elicitation dataset (data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)).
Model $a$ $b$ $\alpha$ $\beta$ SSE RMSE $R^{2}$ AIC BIC
@@
Linear (with intercept) $\alpha+\beta x$
Linear (through origin) cxcx $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
AIC $n\ln(\mathrm{SSE}/n)+2k$, BIC $n\ln(\mathrm{SSE}/n)+k\ln n$; here n=7n{=}7. Parameter counts: k=2k{=}2 for logarithmic, k=2k{=}2 for linear with intercept, and k=1k{=}1 for linear through origin.
Table 3: Goodness-of-fit statistics for Institution 1 minor-error elicitation ($x$ in pages; $y$ in minor-error equivalents).
Comparison.
The logarithmic model substantially improves fit compared to a purely proportional linear rule. Reporting both constrained and unconstrained linear baselines strengthens the robustness of the comparison and avoids conflating functional form with parameter constraints.